A. Find the supplementary angle of 34
A. 146 B. 56 C. 34 D. 68
(2 pts) B. ABC and PQR are supplementary angles and ABC is eleven times as large as PQR. Determine the measure of each angle.

Answer by Earlsdon(6294)

Supplementary angles add up to 180 degrees.
180-34+=+146degrees.
---------------------------
ABC + PQR = 180 degrees. The angles are supplementary.
ABC = 11*(PQR) Substitute this into the first equation.
11*(PQR)+PQR = 180
12*(PQR) = 180 Divide both sides by 12.
PQR = 15 degrees.
ABC = 11*(PQR)
ABC = 11*(15)
ABC = 165 degrees.